OpenStudy (anonymous):

Part A: If (6^2)^x = 1, what is the value of x? Part B: If (6^0)^x = 1, what are the possible values of x?

OpenStudy (anonymous):

just U should know \[(6^{2})^{x}=6^{2x}\]

OpenStudy (fibonaccichick666):

Do you know why what M_A said is true @ImNotLostLikeYou ?

OpenStudy (anonymous):

\[6^{?}=1\]

OpenStudy (anonymous):

i'm struggling with no useful help

OpenStudy (solomonzelman):

a couple of rules you might want to know for this problem: ~ \(\large\color{black}{ (a^{b})^c=a^{~b \times c}}\) ~ \(\large\color{black}{ a^0=1}\) (when \(\large\color{red}{ a \ne 0}\) )

OpenStudy (anonymous):

uh...

OpenStudy (fibonaccichick666):

So, can you tell us what you don't understand? A little more in depth so we know where to start from

OpenStudy (solomonzelman):

your first problem is: \(\large\color{black}{ (6^2)^x=1}\) apply the 1st rule to it.

OpenStudy (anonymous):

the letter times a letter and equals to a letter

OpenStudy (anonymous):

i want to know how to find what the letter means

OpenStudy (fibonaccichick666):

ok, so, a letter, is just a variable. It's a place holder. We use it because it means we can pick ANY numbers and put the number in place of the variable, and the result is always true

OpenStudy (anonymous):

oh ok.

OpenStudy (fibonaccichick666):

Can you please read these few pages, I think they may help you with the background info necessary for this problem http://www.purplemath.com/modules/exponent.htm

OpenStudy (fibonaccichick666):

There are like 5 pages, but the first one should clarify the rules Solomon posted

OpenStudy (anonymous):

I want to find the missing exponent not the missing base

OpenStudy (solomonzelman):

I wouldn't necessarily want to solve the second one, because 1^a=1 , a is all numbers

OpenStudy (solomonzelman):

but we can at least bear with the first problem

OpenStudy (fibonaccichick666):

Those rules are necessary to find x values

OpenStudy (fibonaccichick666):

Which solomon just pointed out, for part two

OpenStudy (fibonaccichick666):

You need to know how to simplify an exponentil expression to an exponent before you can do either problem

OpenStudy (solomonzelman):

If I had: \(\large\color{black}{ (4^3)^c=1}\) then you would rewirte it as: \(\large\color{black}{ 4^{3c}=1}\) So if i had: If I had: \(\large\color{black}{ (6^2)^x=1}\) how would i then write it?

OpenStudy (anonymous):

Part A Answer: x = 0??

OpenStudy (fibonaccichick666):

Please answer Solomon's posed wuestion @ImNotLostLikeYou

OpenStudy (anonymous):

6^2z

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

z is a typo I know

OpenStudy (fibonaccichick666):

so now, you have \[6^{2x}=1\]

OpenStudy (solomonzelman):

then take the log of both sides

OpenStudy (anonymous):

im confused, I have to find the variable of x not z?

OpenStudy (solomonzelman):

and one more rule you need: \(\large\color{black}{ \log(a^b)~~~~\Rightarrow ~~~~b~\log(a)}\)

OpenStudy (solomonzelman):

well, you said that it is z.

OpenStudy (solomonzelman):

I mean that it is x.

OpenStudy (fibonaccichick666):

It's just a variable, a placeholder